The number of grams of a certain radioactive substance present at time t is given by the formula A = 100e^-0.002t, where t is the number of years. Find the number of grams that are present at time t = 0 and time t = 300
Thank you for your help

Answer by stanbon(48568)

The number of grams of a certain radioactive substance present at time t is given by the formula A = 100e^-0.002t, where t is the number of years. Find the number of grams that are present at time t = 0 and time t = 300
-------------------
A(t) = 100e^-0.002t
A(0) = 100*1 = 100 grams
---------
A(300) = 100*e^(-0.002*300) = 100*0.548812
= 54.8812 grams
===================
Cheers,
Stan H.